# Learning Prehensile Dexterity by Imitating and Emulating State-only Observations

### v1

@@ IV Experimental Evaluation
In this section, we report results that i) validate the need for combining imitation with emulation (see Sec. IV-B), ii) demonstrate that CIMER exhibits better sample efficiency compared to baselines and that it can generate more realistic behaviors (see Sec. IV-C), and iii) show that CIMER significantly outperforms baselines, and even the expert in some cases, in terms of both zero-shot generalization and skill transfer to novel objects (see Sec. IV-D). Please refer to the appendices for various additional details.
@@ IV-A Experimental Design
Evaluation Platform: We conducted all our experiments on the widely-used ADROIT Hand [4] – a 30-DoF simulated system (24-DoF hand + 6-DoF floating base) built with MuJoCo Simulator [51].
Data and Expert Policy: We utilized expert DAPG policies [4] to generate 200 state-only observations for each task, comprising both hand and object states. Note that these expert policies were trained using action labels.
Tasks: We consider three prehensile manipulation tasks originally proposed in [4] (see Fig. 3 and Appendix A).
@@
Metrics: We quantify performance in terms of Task success rate (see Appendix B for criteria) and sample efficiency, always reported across five random seeds unless specified.
IV-B Why Care about Emulation?
We first demonstrate why it is not sufficient to merely imitate observations without emulation. Specifically, the drawbacks of merely using a PD controller to track the hand motions generated by the motion generation policy $\Phi$. To this end, we compare CIMER against two baselines: a). Expert Obs. + PD : Observed expert hand motion tracked using a PD controller, and b). Motion Gen. + PD: Reference hand motion (generated by $\Phi$) tracked using a PD controller. We employed the same PD tracking controller with identical PD gains for all three methods and three tasks, and report the task success rate over the 200 initial conditions for each task.
@@ IV-B Why Care about Emulation?
Figure 4: Object tracking is significantly more important than hand tracking for refining hand motions during emulation.
As seen in Table I, CIMER significantly outperforms both Expert Obs. + PD nor Motion Gen. + PD. In fact, CIMER is three to four times more successful. These results support our claim that merely tracking an observed trajectory or imitating expert observations are insufficient to perform prehensile manipulation tasks. This is likely due to the fact that contact effects and grasp forces are ignored and inconsequential when merely trying to replicate an expert’s hand motions. In contrast, introducing emulation allows the robot to “practice” what is learned during imitation, making necessary refinements to reenact the desired object motion.
Another noteworthy observation is that to obtain such significant task improvement in the Door Opening task indeed requires another robot DoF for the motion refinement. Our conjecture is that refining only the robot finger and palm motions (24 DoFs) may not suffice for this particular task. Consequently, for this task, we augment the motion refinement policy by including another DoF of the 6-DoF floating base, which is the reference rotational motion along the hand shaft axis. This adjustment potentially enables the hand to generate larger momentum for effectively turning the handle, thereby leading to further task improvement. The task success rate for this task is 46.5($\pm$3.4)%, if only refining the robot finger and palm motions.
@@
To better understand how emulation improves performance, we qualitatively analyzed the modifications made by the motion refinement policy to compensate for the lack of action labels. As shown in Fig. 5, for the Tool Use task, employing the motion refinement policy results in the hammer hitting the nail closer to its center point, and the force of impact (measured by a touch sensor on the nail head) is also increased. As a result, the nail can be fully driven into the board, leading to improved task performance. For the Object Relocation task, the grasping force applied from each fingertip on the object is noticeably increased, indicating a secure grasp during the initial contact and transportation towards the target position. For the Door Opening task, we observe that the hand initially rotates in one direction, then swiftly changes direction to generate greater momentum for turning the handle (the dashed lines denote the handle turning phase). Therefore, we find that emulation indeed nudges the robot to apply the necessary forces and adjusts the hands’ motion so that the object is manipulated as desired. Finally, it’s important to note that these analyses are only possible due to the hierarchical structure of our framework, in contrast to other end-to-end methods.
IV-C Comparisons to Baselines
@@
No existing framework can directly accommodate all the constraints and challenges of our problem formulation (see Sec. III-A). However, we compare CIMER to the following baselines to demonstrate its relative benefits:
@@
  • Pure RL: We adopt TRPO [52] as a pure RL baseline. This baseline is trained without any offline data, but with per-task reward engineering [4].
@@
  • SOIL: A SOTA state-only imitation method for dexterous manipulation [13] that learns an inverse dynamics model to predict actions from state-only observations. SOIL also requires per-task reward engineering [4].
For SOIL and Pure RL methods, we utilized the same model architectures and training parameters as recommended in their original works, and we include the training parameters for CIMER in Appendix F. For quantify performance in terms of task success rates, with means and standard deviations computed across five random seeds.
@@
As shown in Fig. 6 (top), CIMER exhibits significantly better sample efficiency than SOIL and Pure RL for both Tool Use and Object Relocation tasks. In particular, while the baselines struggle to make any progress on the Object Relocation task, CIMER achieved 100% task success rate. In contrast, baselines were able to achieve higher success rates in the Door Opening task with fewer samples compared to CIMER. This is be partly due to the door opening task likely requiring CIMER to make more significant motion refinements to generate enough momentum that will swing the door open.
@@
A closer examination reveals something important: the baselines tend to exploit the simulator and generate unrealistic or unsafe motions across all tasks. The exemplar rollouts in Fig. 6 (bottom) show that the baselines drive the hand to “throw” the hammer, hit the ball, and penetrate the door handle, while CIMER generates more realistic and stable motions. This is because CIMER’s motion generation policy employs a structured dynamical system that provides a natural and effective motion prior for the motion refinement policy to adjust. These results also suggest that CIMER is better suited for real-world deployment and will not require significant efforts in ensuring safety and efficacy.
IV-D Generalization to new objects
@@
In addition to the objects considered thus far, we evaluated CIMER and the baselines on their ability to generalize to 17 novel objects (3 artifacts and other 14 from the YCB dataset [23]) in the Object Relocation task (see Fig. 7). Note that the Ball is the only default object used for training the expert policy with action labels and for collecting state-only demonstrations.
@@ IV-D1 Zero-shot Generalization
We report the zero-shot generalization performance of each method and the expert across five random seeds in Fig. 8. Given that both the baselines struggled to learn to relocate the default Ball object, it is no surprise that they consistently fail to generalize to novel objects. While the expert achieves a 100% success rate on the default object (ball) with the benefit for action labels, CIMER notably outperforms the expert on the majority of the novel objects. This impressive ability to readily generalize to novel objects could be attributed to CIMER’s hierarchical structure and suggests that CIMER’s motion prior and refinement strategies are robust to changes in the object’s geometric and kinematic properties. Furthermore, CIMER demonstrates the most consistent performance across different training seeds, as indicated by shorter error bars.
Figure 8: CIMER provides impressive zero-shot generalization to novel objects, even outperforming the expert policy in most cases
@@ IV-D2 Skill Transfer to New Objects
We also evaluated if transferring the skills learned on the default object (Ball) to novel objects would improve sample efficiency (see Appendix G). We fine-tuned CIMER’s motion refinement policy on six novel objects, each of which resulted in 50% or lower success rate during zero-shot generalization. We compared against a similarly-finetuned expert policy [4], and two policies trained from scratch (one using CIMER and the other using the expert method).
Figure 9: Skill transfer provides a significant boost to CIMER’s sample efficiency (top two rows), unless the target object is considerably different from the source (bottom row). Solid lines show mean trends and shaded areas show $\pm$ standard deviation over five random seeds.
As shown in Fig. 9, transferring motion refinement skills considerably improves task performance in four of the six objects. On these four objects, CIMER with skill transfer outperforms the baselines trained from scratch as well as the fine-tuned expert. In contrast, skill transfer seems to hurt performance when dealing with other two objects. This could be because skill transfer is counter-productive when target objects are drastically different from the source object, requiring a very different grasping strategy. Once again, CIMER outperforms the expert method on all these novel objects, if trained from scratch or fine-tuned.
@@ V Limitations and Future Work
Our work represents the first task-agnostic and intervention-free approach to learning dexterous prehensile manipulation skills from state-only observations. But it leaves behind several avenues for further improvement. First, given that human videos are ubiquitous online, one could integrate recent advances in motion retargetting [5] with CIMER to directly learning from videos. Second, we do not consider the manipulation of deformable and fragile objects. It is yet unclear how to represent such objects and incentivize emulation for such objects. Third, some objects require a completely different hand trajectory (e.g., grasping a pan by the handle vs. a ball). It might be possible to enable such drastic adaptations by refining the motion of the hand base in addition to the fingers and wrist. Fourth, estimating local contacts using tactile sensors [53, 54, 55] might help refine the hand motion. Last, CIMER is yet to be implemented on physical robots. Leveraging the fact that CIMER outputs PD targets, we plan to deploy on hardware and validate its benefits in the real-world.
Appendix A Task State Space Designs
The state design for each task is as follows:
@@ Appendix A Task State Space Designs
Tool use: For this task, the floating wrist base can only rotate along the $x$ and $y$ axis, so we have $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{26}$. Regarding the object states, unlike the other tasks, where the objects of interest are directly manipulated by the hand, this task requires to modify the environment itself. As a result, except for the hammer positions, orientations and their corresponding velocities $\mathrm{p}^{\text{tool}}_{t},\mathrm{o}^{\text{tool}}_{t},\dot{\mathrm{p}}^{\text{tool}}_{t},\dot{\mathrm{o}}^{\text{tool}}_{t}$ ($\mathbb{R}^{3}$), we also define the nail goal position $\mathrm{p}^{\text{nail}}$ ($\mathbb{R}^{3}$). Finally, we have $\mathrm{o}_{t}=[\mathrm{p}^{\text{tool}}_{t},\mathrm{o}^{\text{tool}}_{t},\dot{\mathrm{p}}^{\text{tool}}_{t},\dot{\mathrm{o}}^{\text{tool}}_{t},\mathrm{p}^{\text{nail}}]\in\mathcal{O}\subseteq\mathbb{R}^{15}$. We use the $\mathrm{p}^{\text{tool}}_{t}$ and $\mathrm{o}^{\text{tool}}_{t}$ as the object tracking states.
@@
Object relocation: For this task, the ADROIT hand is fully actuated, so we have $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{30}$ (24-DoF hand + 6-DoF floating wrist base). Regarding the object states, we define $\mathrm{p}^{\text{target}}$ and $\mathrm{p}^{\text{ball}}_{t}$ as the target and current positions. Then, we compute $\bar{\mathrm{p}}^{\text{ball}}_{t}=\mathrm{p}^{\text{ball}}_{t}-\mathrm{p}^{\text{target}}$, which is the component of $\mathrm{p}^{\text{ball}}_{t}$ in a new coordinate frame that is constructed by $\mathrm{p}^{\text{target}}$ being the origin. We additional include the ball orientation $\mathrm{o}^{\text{ball}}_{t}$ and their corresponding velocities $\dot{\mathrm{p}}^{\text{ball}}_{t}$, $\dot{\mathrm{o}}^{\text{ball}}_{t}$ (all $\mathbb{R}^{3}$). Finally, we have $\mathrm{o}_{t}=[\bar{\mathrm{p}}^{\text{ball}}_{t},\mathrm{o}^{\text{ball}}_{t},\dot{\mathrm{p}}^{\text{ball}}_{t},\dot{\mathrm{o}}^{\text{ball}}_{t}]\in\mathcal{O}\subseteq\mathbb{R}^{12}$. We use the $\bar{\mathrm{p}}^{\text{ball}}_{t}$ as the object tracking states.
@@
For all tasks, the simulation ran at 500HZ. Note that for CIMER and baselines, the robot states remain the same, whereas for object states, we adhere to the design choices proposed in the original works for the baselines [4, 13].
Appendix B Task Success Criteria
We used the same task success criteria as in [4, 13, 44]:
Tool Use: The task is considered successful if at last time step $T$, the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01.
Object Relocation: At each time step $t$, if $\sqrt{|\mathrm{p}^{\text{target}}-\mathrm{p}^{\text{ball}}_{t}|^{2}}<0.10$, then we have $\rho(t)=1$. The task is considered successful if $\sum_{t=1}^{T}\rho(t)>10$.
Door Opening: The task is considered successful if at last time step $T$, the door opening angle is larger than 1.35 rad.
Learning Prehensile Dexterity by Imitating and Emulating State-only Observations
Yunhai Han, Zhenyang Chen, Harish Ravichandar
Georgia Institute of Technology
Email: {yhan389,zchen927,harish.ravichandar}@gatech.edu
Appendix C CIMER Pseudo-code

### v3

@@ IV Experimental Evaluation
Our experiments pose the following questions: i) Is emulation necessary? (Sec. IV-B), ii) Is CIMER more sample efficient than baselines? (Sec. IV-C), iii) Does CIMER generate robust and realistic motions? (Sec. IV-C), and iv) Can CIMER generalize and adapt to novel objects (Sec. IV-D).
@@ IV-A Experimental Design
Evaluation Platform: We used the widely-used ADROIT Hand [4] – a 30-DoF simulated system (24-DoF hand + 6-DoF floating base) built with MuJoCo Simulator [53].
Data and Expert Policy: We utilized expert DAPG policies [4] trained using action labels to generate 200 state-only observations (both hand and object states) for each task.
Tasks: We consider three widely-used prehensile tasks originally proposed in [4] (see Fig. 3 and Appendix A).
@@
To ensure fairness and reproducibility, we adopt the same success criteria originally proposed in [4].
Metrics: We quantify performance in terms of Task success rate (see [4, 13, 45] for criteria) and sample efficiency, reported across five random seeds unless specified otherwise.
Figure 4: Emulating the observed object motion is significantly more effective than emulating the observed hand motion.
IV-B Need for Emulation
We first demonstrate why it insufficient to imitate state-only observations without emulation. To this end, we compare CIMER against three baselines that solely rely on imitation: a). Expert Obs. [4] + PD: Reference hand motion generated by the expert policy trained with action labels, b). Motion Gen. + PD: Reference hand motion generated by $\Phi$, and c). DexTransfer [12] + PD: Reference hand motion generated by a policy trained on an augmented dataset (generated by simulating perturbed observations and augmenting new trajectories deemed successful by task-specific success criteria). We used the same PD controller with identical gains for all tasks and methods, and report success rates over 200 initial conditions.
@@ IV-B Need for Emulation
As seen in Table I, CIMER significantly outperforms all three baselines. These results support our claim that merely tracking an observed expert trajectory is insufficient for prehensile dexterous manipulation since contact effects and grasp forces are ignored when merely imitating state-only observations. Though DexTransfer uses the simulator to evaluate the effectiveness of perturbed trajectories, it only learns from successful trajectories and disregards ones that fail. In contrast, CIMER uses emulation to “practice” what it learned during imitation, refining the hand motion to reproduce the observed object motion. We also observed that refining only the robot finger and palm motions may not always be sufficient. Specifically, we found that doing so improves task success to only 46.5($\pm$3.4)in the Door Opening task. However, the success rate jumps to 76.3($\pm$13.4)when CIMER also refines the motion of the 6-DoF hand base. This more comprehensive refinement is likely helping the hand generate larger momentum to successfully turn the door handle.
@@
To better understand how emulation improves performance, we qualitatively analyzed the modifications made by the motion refinement policy to compensate for the lack of action labels in each task (see Fig. 5). Tool Use: The hammer hits the nail closer to its center point and with greater force (measured by a touch sensor on the nail head), helping drive the nail completely into the board. Object Relocation: the grasping force applied from each fingertip on the object is noticeably increased, producing a more secure grasp during the initial contact and transportation towards the target. Door Opening: the hand initially rotates in one direction, then swiftly changes direction to generate greater momentum for turning the handle (the dashed lines denote the handle turning phase). These observations suggest that emulation indeed nudges the robot to apply the necessary forces and adjusts the hand’s motion so that the object is manipulated as desired.
IV-C Sample Efficiency and Realism
@@
Though no existing methods can accommodate all the constraints and challenges of our problem setting (see Sec. III-A), we compare CIMER to the following baselines to evaluate its relative benefits in terms of sample efficiency and realism:
@@
  • Pure RL: We trained TRPO [54] as a pure RL baseline without offline data, but with task-specific reward shaping [4] to evaluate the benefits of leveraging observations.
@@
  • SOIL [13]: We also evaluated CIMER against this SOTA state-only imitation method for dexterous manipulation, which also leverages both expert observations and interactions with a simulator. Unlike CIMER, SOIL also requires task-specific reward engineering [4].
For both baselines, we utilized the same model architectures and training parameters as recommended in their original works. We report all results over five random seeds.
@@
As shown in Fig. 6 (top), CIMER exhibits significantly better sample efficiency than SOIL and Pure RL for both Tool Use and Object Relocation tasks. In particular, while the baselines struggle to make any progress on the Object Relocation task, CIMER achieved 100% task success rate. This is likely the baselines end up learning a fragile strategy of bouncing the ball against the table instead of grasping it (see Fig. 6 (bottom)). In contrast, baselines were able to achieve higher success rates in the Door Opening task with fewer samples compared to CIMER. This is partly due to the door opening task likely requiring more significant motion refinements to generate enough momentum that will swing the door open. Further, unlike the baselines, CIMER does not benefit from task-specific reward design.
@@
A closer examination revealed something important: the baselines tend to exploit the simulator and generate unrealistic or aggresive motions across all tasks. Exemplar rollouts in Fig. 6 (bottom) show that the baselines tend to drive the hand to “throw” the hammer, hit and bounce the ball, and penetrate the door handle, while CIMER generates more realistic and stable motions. This is because CIMER’s motion generation policy employs a structured dynamical system that provides a realistic and effective motion prior for the motion refinement policy to adjust. These results also suggest that CIMER is better suited for real-world deployment and will not likely require significant efforts in ensuring safety and efficacy.
To quantitatively evaluate realism and the suitability for real-world implementation, we evaluated the robustness of each method to systematic changes in all mass and damping coefficients (see Appendix C for details). The results reveal that CIMER results in comparable or better robustness, even outperforming the expert policy trained using action labels on the Relocation task. Though Pure RL and SOIL seem to be more robust than CIMER on the Door task, note that qualitative analysis revealed that both Pure RL and SOIL tend to exploit the simulator and generate unrealistic and agrressive motions (see Fig. 6 (bottom)).
IV-D Generalization and Adaptation to Novel Objects
@@
In addition to the objects considered thus far, we evaluated CIMER and the baselines on their ability to generalize to 17 novel objects (3 synthetic objects and 14 objects from the YCB dataset [24]) in the Object Relocation task (see Fig. 8). Note that the Ball is the only default object used for training the expert policy and for collecting state-only demonstrations.
Figure 7: The evaluation of the robustness of each method to systematic changes in mass and damping coefficients.
@@ IV-D1 Zero-shot Generalization
We report the zero-shot generalization performance of each method and the expert across five random seeds in Fig. 9. Given that both the baselines struggled to learn to relocate the default Ball object, it is no surprise that they consistently fail to generalize to novel objects. While the expert achieves a 100% success rate on the default object (ball) with the benefit for action labels, CIMER notably outperforms the expert on the majority of the novel objects. This impressive ability to readily generalize to novel objects could be attributed to CIMER’s hierarchical structure and suggests that CIMER’s motion prior and refinement strategy are somewhat robust to changes in the object’s geometric and kinematic properties. Further, CIMER demonstrates the most consistent performance across different training seeds, as indicated by shorter error bars.
Figure 9: CIMER provides impressive zero-shot generalization to novel objects, even outperforming the expert policy in most cases.
@@ IV-D2 Skill Transfer to New Objects
We also evaluated if transferring the skills learned on the default object (Ball) to novel objects would improve sample efficiency (see Appendix G for experiments on learning from scratch). Freezing the motion generation policy, we fine-tuned CIMER’s motion refinement policy on six novel objects (each of which resulted in 50% or lower success rate during zero-shot generalization). We compared against a similarly-finetuned expert policy [4], and two additional policies trained from scratch (one using CIMER and the other using the expert method).
Figure 10: Skill transfer significantly boosts CIMER’s sample efficiency (top two rows), unless the target object is considerably different from the source (bottom row). Solid lines show mean trends and shaded areas show $\pm$ standard deviation over five random seeds.
As shown in Fig. 10, transferring motion refinement skills considerably improves sample efficiency in four of the six objects. On these four objects, CIMER with skill transfer outperforms the baselines trained from scratch as well as the fine-tuned expert. In contrast, skill transfer hurts sample efficiency when dealing with other two objects. This is likely because skill transfer is counter-productive when target objects are drastically different from the source object, requiring a significantly different grasping strategy. In such cases, CIMER has to “unlearn” the previous skill before adapting to the novel object.
@@ V Limitations and Future Work
Our work represents the first task-agnostic and intervention-free approach to learning dexterous prehensile manipulation skills from state-only observations. But it still leaves behind several avenues for further improvement. First, given that human videos are ubiquitous online, one could integrate recent advances in motion retargetting [5] with CIMER to directly learning from videos. Second, we do not consider the manipulation of deformable and fragile objects. It is yet unclear how to represent such objects in way that makes incentivizing emulation easier and tractable. Third, some objects and contexts require significantly different strategies (e.g., grasping a pencil to write vs. a ball to throw). It might be possible to enable such drastic adaptations by refining all available degrees of freedom of the hand and accounting for affordances. Fourth, reasoning about local contact events using tactile sensors [55, 56, 57] might help improve refinement and tackle contact-rich tasks (e.g., peg-in-hole). Last, while our robustness experiments show promise, CIMER is yet to be implemented on physical robots. Leveraging the fact that it generates PD targets, we plan to deploy CIMER on hardware and validate its benefits.
Appendix A State Space Designs
Our simulation ran at 500 Hz and the state design for each task is as follows:
@@ Appendix A State Space Designs
Tool use: The floating hand base can only rotate along the $x$ and $y$ axes, resulting in $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{26}$. Unlike other tasks, where the objects of interest are directly manipulated by the hand, this task requires indirect manipulation of the nail. As such, in addition to the hammer position, orientation, and corresponding velocities $\mathrm{p}^{\text{tool}}_{t},\mathrm{o}^{\text{tool}}_{t},\dot{\mathrm{p}}^{\text{tool}}_{t},\dot{\mathrm{o}}^{\text{tool}}_{t}$ ($\mathbb{R}^{3}$), we define the nail goal position $\mathrm{p}^{\text{nail}}$ ($\mathbb{R}^{3}$). Finally, we have $\mathrm{o}_{t}=[\mathrm{p}^{\text{tool}}_{t},\mathrm{o}^{\text{tool}}_{t},\dot{\mathrm{p}}^{\text{tool}}_{t},\dot{\mathrm{o}}^{\text{tool}}_{t},\mathrm{p}^{\text{nail}}]\in\mathcal{O}\subseteq\mathbb{R}^{15}$. We use the $\mathrm{p}^{\text{tool}}_{t}$ and $\mathrm{o}^{\text{tool}}_{t}$ as the object tracking states.
@@
Object relocation: The ADROIT hand is fully actuated, so we have $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{30}$ (24-DoF hand + 6-DoF floating wrist base). Regarding the object states, we define $\mathrm{p}^{\text{target}}$ and $\mathrm{p}^{\text{ball}}_{t}$ as the target and current positions. Then, we compute $\bar{\mathrm{p}}^{\text{ball}}_{t}=\mathrm{p}^{\text{ball}}_{t}-\mathrm{p}^{\text{target}}$, which is the component of $\mathrm{p}^{\text{ball}}_{t}$ in a new coordinate frame that is constructed by $\mathrm{p}^{\text{target}}$ being the origin. We additional include the ball orientation $\mathrm{o}^{\text{ball}}_{t}$ and their corresponding velocities $\dot{\mathrm{p}}^{\text{ball}}_{t}$, $\dot{\mathrm{o}}^{\text{ball}}_{t}$ (all $\mathbb{R}^{3}$). Finally, we have $\mathrm{o}_{t}=[\bar{\mathrm{p}}^{\text{ball}}_{t},\mathrm{o}^{\text{ball}}_{t},\dot{\mathrm{p}}^{\text{ball}}_{t},\dot{\mathrm{o}}^{\text{ball}}_{t}]\in\mathcal{O}\subseteq\mathbb{R}^{12}$. We use the $\bar{\mathrm{p}}^{\text{ball}}_{t}$ as the object tracking states.
@@
Appendix B CIMER Hyperparameters
TABLE II: Training details
MLP Hidden Layer RL Learning Rate GAEγ
$(256,128)$ 2e-6 0.98
PPO clip threshold PPO epochs GAEλ
0.2 8 0.97
In addition, we used $k^{h}=5,k^{o}=5,r_{t}^{b}=25$ in all tasks to compute the tracking rewards introduced in Sec. III-D.
Appendix C Robustness Experiments
For the robustness experiments in Sec. IV-C, we varied mass and damping since they are an important subset of differences between simulation and hardware. Specifically, we varied all mass and damping coefficients along eleven equally spaced evaluation points at $[0\%,10\%,20\%,...,100\%]$ variation. For each variation, we sampled mass and damping coefficients from a uniform distribution, with the lower and upper bounds being multiples of the default values. The lower and upper bound values ($l_{j}$ and $u_{j}$) for each variation case ($j$) are
$\forall j\in\{0,0.1,0.2,...,1.0\},l_{j}=1-0.8j,u_{j}=1+4.0j.$
After sampling a particular combination of mass and damping parameters, we tested each policy for 200 episodes with varying goal locations. We repeated this process 20 times for each variation, and report the resulting average success rates.
Appendix D CIMER Pseudo-code